# Retrieval-Augmented Generation-based Relation Extraction

### v1

@@ 3 Methodology
The rest of the section explains the details of how each module of our proposed approach at Figure 2 works under specific subsections.
@@ 3.1 Retrieval
A user submits a sentence (query) along with a pair of entities (head and tail entities) that might have a relation to the Retrieval module. Then, the Retriever sends this query to the Data Augmentation module, which extends the original query with a semantically similar sentence from training dataset, as an example given in Figure 4. “Result Refinement” in this module applies post-processing approaches when results are returned from the Generation module. The “Result Refinement” consists of a couple of response processing steps, such as refining prefixes (e.g., changing “founded” to “org:founded”), and converting “no relation” answers into “no_relation” as defined in the predefined relation types. This ensures that only relation types are returned instead of explanations from the LLMs.
@@ 3.2 Data Augmentation
The Data Augmentation module contains an Embedding Database (DB) consisting of embeddings of training data, computed using the Sentence BERT (SBERT) model [26]. We used the “all-MiniLM-L6-v2”  [^2] The model is available at https://huggingface.co/sentence-transformers/all-MiniLM-L12-v2 version of SBERT in our approach. Additionally, the query embedding is computed by SBERT within this module. Subsequently, the system calculates similarity scores between the embeddings of each line of training sentences in the Embedding DB and the query embeddings, utilizing the cosine similarity metric. The system considers the most similar embeddings from the Embedding DB as an example of the query. Both the query and the similar sentence example are then inputted into the prompt generator, which prepares prompts by considering the template defined in Figure 3. In other words, prompt generator regenerates the user query along with the relevant sentence that was founded at the Embedding DB. One of the generated prompts can be seen in Figure 4. The prompt generator then inputs the prompt (regenerated user query), consisting of the query and its similar example, into the Generation module.
@@ 4 Evaluation
In this section, we examine performance of our RAG4RE work. We first introduce experimental settings in Section 4.1 and then give results of our experiments in Section 4.2. Afterwards, false predictions are analyzed in Section 4.3. Lastly, we discuss the performance of our approach with previous works.
@@ 4.1 Experimental Setup
In our work, we assess the effectiveness of our RAG4RE approach by leveraging well-established RE benchmarks, including TACRED [11], TACREV [12], Re-TACRED [13], and SemEval [14] RE datasets. These benchmarks include head and tail entities in given sentences, as well as ground truth relation types between these entities. These widely recognized datasets serve as invaluable resources for assessing the efficiency and performance of our approach. Further insights into the dataset can be seen below, and in Table 1.
@@
Additionally, we evaluate our proposed method on these datasets by integrating various LLMs, such as Flan T5 (XL and XXL), Llama2-7b, and Mistral-7B-Instruct-v0.2 models. The prompt template for our RAG4RE includes a relevant sentence provided by the user (See the prompt template in Figure 3), whereas the simple query does not include any relevant sentence regarding the query. This comprehensive evaluation helps us understand the performance of our approach across various LLMs. Lastly, we compare our RAG4RE approach with simple query (Vanilla LLM prompting) in terms of micro F1, Recall, and Precision scores, since these benchmark datasets are imbalance. To compute these metrics, we leverage metrics library of sklearn  [^7] Sklearn evaluation library:  https://scikit-learn.org/stable/modules/model_evaluation.html.
@@ 4.2 Results
Our experiments were conducted using the four benchmark datasets mentioned above. Firstly, we assessed the performance of our proposed RAG4RE and compared it to that of a simple query (sentence), Vanilla LLM prompting, in terms of micro F1 score. As mentioned earlier, our evaluation criteria took into account the micro F1 score, Recall, and Precision due to the imbalanced labelling of the datasets. Furthermore, we explored how our approach enhances the performance of LLM responses. This was achieved by incorporating a similar example sentence into the prompt template alongside the query sentence in our proposed RAG4RE approach (See Data Augmentation at Figure 2). We compared the outcomes of a simple query without any relevant sentence to our results (See Table 2 and Table 3).
We leveraged various LLMs to conduct our experiments and evaluated the performance of our proposed RAG4RE. Our RAG4RE has achieved remarkable results when compared its performance to that of the simple query at Table 2 and Figure 5. It can be seen that our RAG4RE outperforms the simple query sending to the LLMs in TACRED, TACREV, and Re-TACRED despite the fact that the language model has been changed. The highest F1 scores amongst the results at Table 2 have been accomplished 86.6%, 88.3% and 73.3% of F1 scores on the TACRED, TACREV and Re-TACRED, respectively. These outstanding scores were achieved through the integration of the Flan T5-XL model into the Generation module. Unfortunately, our RAG4RE fell short of achieving comparable performance on the SemEval dataset. This is primarily because the predefined relations (target relation labels) within this dataset cannot be directly extracted from the sentence tokens themselves.
The notable improvement observed in RAG4RE’s results can be primarily attributed to the incorporation of a relevant example sentences extracted from benchmark datasets’ training sentences into the user query. As pointed out in [7], RAG operates akin to an open book exam, and adding the relevant sentence regarding a query sentence plays a facilitating role in understanding of the query by the LLMs in the Generator module of our approach at Figure 2. This understanding is supported by the results outlined in Table 2.
@@
Flan T5-XL simple query $91.091.0$ $79.279.2$ $84.784.7$ $97.297.2$ $49.049.0$ $65.165.1$ $69.569.5$ $73.073.0$ $71.271.2$ $20.4120.41$ $12.3212.32$ $15.3715.37$
RAG4RE $84.584.5$ $88.888.8$ 86.6 $84.584.5$ $92.492.4$ 88.3 $63.063.0$ $87.787.7$ 73.3 $17.1617.16$ $11.9311.93$ $14.0714.07$
Flan T5-XXL simple query $94.994.9$ $44.744.7$ $60.960.9$ $98.298.2$ $27.127.1$ $42.442.4$ $70.270.2$ $39.339.3$ $50.450.4$ 13.64 13.4 13.52
RAG4RE $93.393.3$ $61.061.0$ $73.873.8$ $91.791.7$ $82.382.3$ $86.786.7$ $78.078.0$ $53.653.6$ $63.563.5$ 17.32 15.39 16.29
@@
Consequently, it is clear that our RAG4RE approach, consisting of the relevant example sentence about the query, has improved F1 scores on benchmark datasets whose predefined relations (target relationship labels) are coming from the given sentence tokens e.g., TACRED, TACREV and Re-TACRED, when compared its results to those of a simple query (Vanilla LLM prompting) as demonstrated in Table 2 and Figure 5.
@@ 4.3 Error Analysis
Decoder only LLMs, such as Mistral and Llama2, are prone to producing hallucinatory results when a simple prompt is sent to those models [18]. In the responses of the experiments conducted with Mistral-7B-Instruct-v0.2 and Llama2-7b, we observed such relation types which have not been defined in the relation types of the prompt template (See the example prompt in Figure 3). We analyze False Negative and False Positive relation predictions of three language model types in Table 3 and Figure 6.
@@
Flan T5-XL simple query $956956$ $37263726$ $316316$ $81808180$ $24812481$ $36763676$ $13061306$ $10761076$
RAG4RE $19901990$ $19871987$ $23372337$ $13901390$ $40134013$ $17351735$ $15641564$ $829829$
Llama2-7b simple query 26 15336 18 15438 18 13327 $0$ $25792579$
RAG4RE 1549 7258 1184 $76$ 212 12937 $0$ $26032603$
Mistral-7B-Instruct-v0.2 simple query 82 13970 52 14830 233 12759 $0$ $20552055$
RAG4RE 509 11333 200 12557 1546 9535 $0$ $20992099$
@@ 4.4 Discussion
In our experiments, we compared two methods: (i) RAG-based Relation Extraction (RAG4RE) and (ii) a simple query (Vanilla LLM prompting) which lacks the inclusion of the relevant example sentence described in Figure 3. Our findings indicate a notable enhancement in F1 scores when employing the RAG4RE approach over the simple query method on the TACRED dataset and its variations, as outlined in Table 2. This improvement stems from the integration of a relevant sentence into the prompt template, as illustrated in Figure 3. This inclusion of relevant example sentence facilitates the predictions made by the LLM in the Generation module of our proposed architecture, as depicted in Figure 2.
The incorporation of this relevant sentence serves to mitigate hallucinations in the LLM’s responses, subsequently reducing the occurrence of false predictions, as demonstrated in Table 3 and  Figure 6. Our assessment of the RAG4RE approach’s effectiveness is based on the integration of Flan T5-XL into the LLM within Figure 2, given that our approach, combined with Flan T5-XL, yields the highest F1 scores across benchmarks except for the SemEval (See Table 2).
We present an analysis of the performance of our RAG4RE approach, comparing its F1 score with both LLM-based methods and State-of-the-Art (SoTA) Relation Extraction (RE) techniques reported in current literature. In terms of LLM-based RE approaches, our RAG4RE consistently outperforms other methods utilizing LLMs, as illustrated in Table 5, across all benchmark datasets except for SemEval. The reason for the superior performance of our RAG4RE, as depicted in Table 5, is largely attributed to the absence of relevant sentence addition in the prompt templates of both LLMQA4RE [22] and RationaleCL [27]. Notably, both competing methods did not incorporate relevant sentences into their prompt templates.
Furthermore, we conducted a comparison with works achieving the best results on these benchmarks in the literature. Remarkably, our RAG4RE surpasses the State-of-the-Art results on both TACRED and TACREV datasets, achieving F1 scores of 86.8% and 88.3% respectively, as illustrated in Table 4.

### v2

@@ 3 Methodology
The rest of the section explains the details of how each module of our proposed approach at Figure˜2 works under specific subsections.
@@ 3.1 Retrieval
A user submits a sentence (query) along with a pair of entities (head and tail entities) that might have a relation to the Retrieval module. Then, the Retriever sends this query to the Data Augmentation module, which extends the original query with a semantically similar sentence from training dataset, as an example given in Figure˜4. “Result Refinement” in this module applies post-processing approaches when results are returned from the Generation module. The “Result Refinement” consists of a couple of response processing steps, such as refining prefixes (e.g., changing “founded” to “org:founded”), and converting “no relation” answers into “no_relation” as defined in the predefined relation types. This ensures that only relation types are returned instead of explanations from the LLMs.
@@ 3.2 Data Augmentation
The Data Augmentation module contains an Embedding Database (DB) consisting of embeddings of training data, computed using the Sentence BERT (SBERT) model [26]. We used the “all-MiniLM-L6-v2”  [^2] The model is available at https://huggingface.co/sentence-transformers/all-MiniLM-L12-v2 version of SBERT in our approach. Additionally, the query embedding is computed by SBERT within this module. Subsequently, the system calculates similarity scores between the embeddings of each line of training sentences in the Embedding DB and the query embeddings, utilizing the cosine similarity metric. The system considers the most similar embeddings from the Embedding DB as an example of the query. Both the query and the similar sentence example are then inputted into the prompt generator, which prepares prompts by considering the template defined in Figure˜3. In other words, prompt generator regenerates the user query along with the relevant sentence that was founded at the Embedding DB. One of the generated prompts can be seen in Figure˜4. The prompt generator then inputs the prompt (regenerated user query), consisting of the query and its similar example, into the Generation module.
@@ 4 Evaluation
In this section, we examine performance of our RAG4RE work. We first introduce experimental settings in Section˜4.1 and then give results of our experiments in Section˜4.2. Afterwards, false predictions are analyzed in Section˜4.3. Lastly, we discuss the performance of our approach with previous works.
@@ 4.1 Experimental Setup
In our work, we assess the effectiveness of our RAG4RE approach by leveraging well-established RE benchmarks, including TACRED [11], TACREV [12], Re-TACRED [13], and SemEval [14] RE datasets. These benchmarks include head and tail entities in given sentences, as well as ground truth relation types between these entities. These widely recognized datasets serve as invaluable resources for assessing the efficiency and performance of our approach. Further insights into the dataset can be seen below, and in Table˜1.
@@
Additionally, we evaluate our proposed method on these datasets by integrating various LLMs, such as Flan T5 (XL and XXL), Llama2-7b, and Mistral-7B-Instruct-v0.2 models. The prompt template for our RAG4RE includes a relevant sentence provided by the user (See the prompt template in Figure˜3), whereas the simple query does not include any relevant sentence regarding the query. This comprehensive evaluation helps us understand the performance of our approach across various LLMs. Lastly, we compare our RAG4RE approach with simple query (Vanilla LLM prompting) in terms of micro F1, Recall, and Precision scores, since these benchmark datasets are imbalance. To compute these metrics, we leverage metrics library of sklearn  [^7] Sklearn evaluation library:  https://scikit-learn.org/stable/modules/model_evaluation.html.
@@ 4.2 Results
Our experiments were conducted using the four benchmark datasets mentioned above. Firstly, we assessed the performance of our proposed RAG4RE and compared it to that of a simple query (sentence), Vanilla LLM prompting, in terms of micro F1 score. As mentioned earlier, our evaluation criteria took into account the micro F1 score, Recall, and Precision due to the imbalanced labelling of the datasets. Furthermore, we explored how our approach enhances the performance of LLM responses. This was achieved by incorporating a similar example sentence into the prompt template alongside the query sentence in our proposed RAG4RE approach (See Data Augmentation at Figure˜2). We compared the outcomes of a simple query without any relevant sentence to our results (See Table˜2 and Table˜3).
We leveraged various LLMs to conduct our experiments and evaluated the performance of our proposed RAG4RE. Our RAG4RE has achieved remarkable results when compared its performance to that of the simple query at Table˜2 and Figure˜5. It can be seen that our RAG4RE outperforms the simple query sending to the LLMs in TACRED, TACREV, and Re-TACRED despite the fact that the language model has been changed. The highest F1 scores amongst the results at Table˜2 have been accomplished 86.6%, 88.3% and 73.3% of F1 scores on the TACRED, TACREV and Re-TACRED, respectively. These outstanding scores were achieved through the integration of the Flan T5-XL model into the Generation module. Unfortunately, our RAG4RE fell short of achieving comparable performance on the SemEval dataset. This is primarily because the predefined relations (target relation labels) within this dataset cannot be directly extracted from the sentence tokens themselves.
The notable improvement observed in RAG4RE’s results can be primarily attributed to the incorporation of a relevant example sentences extracted from benchmark datasets’ training sentences into the user query. As pointed out in [7], RAG operates akin to an open book exam, and adding the relevant sentence regarding a query sentence plays a facilitating role in understanding of the query by the LLMs in the Generator module of our approach at Figure˜2. This understanding is supported by the results outlined in Table˜2.
@@
Flan T5-XL simple query $91.0$ $79.2$ $84.7$ $97.2$ $49.0$ $65.1$ $69.5$ $73.0$ $71.2$ $20.41$ $12.32$ $15.37$
RAG4RE $84.5$ $88.8$ 86.6 $84.5$ $92.4$ 88.3 $63.0$ $87.7$ 73.3 $17.16$ $11.93$ $14.07$
Flan T5-XXL simple query $94.9$ $44.7$ $60.9$ $98.2$ $27.1$ $42.4$ $70.2$ $39.3$ $50.4$ 13.64 13.4 13.52
RAG4RE $93.3$ $61.0$ $73.8$ $91.7$ $82.3$ $86.7$ $78.0$ $53.6$ $63.5$ 17.32 15.39 16.29
@@
Consequently, it is clear that our RAG4RE approach, consisting of the relevant example sentence about the query, has improved F1 scores on benchmark datasets whose predefined relations (target relationship labels) are coming from the given sentence tokens e.g., TACRED, TACREV and Re-TACRED, when compared its results to those of a simple query (Vanilla LLM prompting) as demonstrated in Table˜2 and Figure˜5.
@@ 4.3 Error Analysis
Decoder only LLMs, such as Mistral and Llama2, are prone to producing hallucinatory results when a simple prompt is sent to those models [18]. In the responses of the experiments conducted with Mistral-7B-Instruct-v0.2 and Llama2-7b, we observed such relation types which have not been defined in the relation types of the prompt template (See the example prompt in Figure˜3). We analyze False Negative and False Positive relation predictions of three language model types in Table˜3 and Figure˜6.
@@
Flan T5-XL simple query $956$ $3726$ $316$ $8180$ $2481$ $3676$ $1306$ $1076$
RAG4RE $1990$ $1987$ $2337$ $1390$ $4013$ $1735$ $1564$ $829$
Llama2-7b simple query 26 15336 18 15438 18 13327 0 $2579$
RAG4RE 1549 7258 1184 $76$ 212 12937 0 $2603$
Mistral-7B-Instruct-v0.2 simple query 82 13970 52 14830 233 12759 0 $2055$
RAG4RE 509 11333 200 12557 1546 9535 0 $2099$
@@ 4.4 Discussion
In our experiments, we compared two methods: (i) RAG-based Relation Extraction (RAG4RE) and (ii) a simple query (Vanilla LLM prompting) which lacks the inclusion of the relevant example sentence described in Figure˜3. Our findings indicate a notable enhancement in F1 scores when employing the RAG4RE approach over the simple query method on the TACRED dataset and its variations, as outlined in Table˜2. This improvement stems from the integration of a relevant sentence into the prompt template, as illustrated in Figure˜3. This inclusion of relevant example sentence facilitates the predictions made by the LLM in the Generation module of our proposed architecture, as depicted in Figure˜2.
The incorporation of this relevant sentence serves to mitigate hallucinations in the LLM’s responses, subsequently reducing the occurrence of false predictions, as demonstrated in Table˜3 and  Figure˜6. Our assessment of the RAG4RE approach’s effectiveness is based on the integration of Flan T5-XL into the LLM within Figure˜2, given that our approach, combined with Flan T5-XL, yields the highest F1 scores across benchmarks except for the SemEval (See Table˜2).
We present an analysis of the performance of our RAG4RE approach, comparing its F1 score with both LLM-based methods and State-of-the-Art (SoTA) Relation Extraction (RE) techniques reported in current literature. In terms of LLM-based RE approaches, our RAG4RE consistently outperforms other methods utilizing LLMs, as illustrated in Table˜5, across all benchmark datasets except for SemEval. The reason for the superior performance of our RAG4RE, as depicted in Table˜5, is largely attributed to the absence of relevant sentence addition in the prompt templates of both LLMQA4RE [22] and RationaleCL [27]. Notably, both competing methods did not incorporate relevant sentences into their prompt templates.
Furthermore, we conducted a comparison with works achieving the best results on these benchmarks in the literature. Remarkably, our RAG4RE surpasses the State-of-the-Art results on both TACRED and TACREV datasets, achieving F1 scores of 86.8% and 88.3% respectively, as illustrated in Table˜4.